# Estimating Spatially Resolved Radiation Fields Using Neural Networks

### v1

@@ 2.2 Data normalization
Figure 1: Distribution of relative voxel fluences $\Phi^{3}_{\mathrm{\gamma}}$ inside datasets DS-01 and DS-02 starting from 0.0 on the X-axis.
Figure 2: Distribution of relative voxel fluences $\Phi^{3}_{\mathrm{\gamma}}$ inside datasets DS-01 and DS-02 starting from $10^{-1}$ on the X-axis.
Figure 3: Distribution of relative voxel fluences $\Phi^{3}_{\mathrm{\gamma}}$ inside dataset DS-03 starting from 0.0 on the X-axis.
Figure 4: Distribution of relative voxel fluences $\Phi^{3}_{\mathrm{\gamma}}$ inside dataset DS-03 starting from $10^{-1}$ on the X-axis.
In principle, radiation fields inside the dataset are already normalized, as the photon spectra are normalized to probability distributions $p(E_{\mathrm{\gamma}})$ and the fluences are normalized per single primary photon from the simulation. However, due to the numerical nature of the datasets, after loading a single radiation field from a dataset, we performed an extra normalization step for the data to limit the occurring gradients and avoid numerical instability as well as introducing bias because of high total values. Thus this extra normalization step effectively mitigates the severe class imbalance in all our dataset towards low fluence voxels. This has been visualized in the global fluence histograms in Fig. 4 for DS-01 and DS-02 and in Fig. 4 for DS-03. The histograms were generated using the MaxNorm function to transform the fluences to $[0,1]$. From both figures we can draw, that the majority of voxels has a fluence lower then $5\,\%$ of the maximum fluence, which makes sense, as the major part of the radiation field is indeed scatter radiation. However, this imbalance could potentially lead to a misreconstruction of the primary beam. Especially for DS-03 from Fig. 4, we can see that the class imbalance is also present inside the primary beam component. Not as severely expressed, datasets DS-01 and DS-02 exhibit an unequal distribution of fluences as well, as visible from Fig. 4. Therefore, we examine the influence of different normalization functions regarding the reconstruction capabilities, to find the best fit to reconstruct the high and low fluence regions with sufficient results.
@@
The idea of using alternative normalizations in the interval of $[0,1]$ is to create a more equal distribution of gradient strengths over the range of values to reconstruct. We pose it as a hypothesis, that this is of importance for our use case, as the dosimetry is regarding relative instead of absolute errors per voxel. Thus, it can be beneficial to elevate differences between small fluences and flatten the differences between higher fluences to give small deviations on small scales more impact. A classical solution for this issue is to use a transformation to the log-space, which is in general a reasonable choice for the following training of neural networks. Unfortunately, due to numerical instabilities, the steep ascend of the logarithm at very small values can become unstable under inversion and thus break the bijective property required for normalization. This instability can be observed for fluences of a magnitude lower than $10^{-5}$. As it can be seen from Tab. 1, for DS-03 the interval of $10^{-7},10^{-8}]$ is of interest for the reconstruction as well. To mitigate this issue we added a scaling factor $\alpha$ prior to the log-transformation. We test two values for $\alpha$: $10^{3}$ and $10^{5}$. As an alternative way to encounter the high dynamic range, we doubled the range of our previously linearly spread out normalization from $\Phi^{3}_{\mathrm{\gamma}}\in[0,1]$ to $\Phi^{3}_{\mathrm{\gamma}}\in[-1,1]$ which additionally separates the low fluence voxels from the few high energy voxels by the sign. We distinguish between those two variants of the MaxNorm by indicating them as $\mathrm{MaxNorm}_{0}^{1}$ and $\mathrm{MaxNorm}_{-1}^{+1}$.
@@ 2.3 Architectures
In order to access the effectivity of our neural network architectures, we implemented each architectures neural network model in three variants to handle the different provided parameters for each dataset. With this approach, for each architecture there is one model that can only handle X-ray tube rotations, one that can additionally handle varying spectra and one that can additionally handle varying tube distances.
In general, our models are all estimating the relative topology of the spatial radiation field distribution conserving the relation of intensities between all voxels inside one specific radiation field. This is important, as the actual fluence scales between different settings of the beam would introduce a bias towards high X-ray tube currents and thus harm the training of the neural networks. On the downside, in the real world application, this inherently requires our approach to run concurrently to at least one area dosimeter that reports doserates at a known point in space. As those devices, especially those with real-time measurement capabilities, are known to exhibit a certain energy and thus spectrum dependence, one needs to correct those devices accordingly. This is specifically of importance in our targeted use case, as we aim to predict the scatter field around the patient, where the spectrum is not homogeneous and difficultly predictable. Thus, we decided to train all our networks not of the actual target measurand, e.g., air kerma rate $\dot{K}_{\mathrm{air}}$ or $\mathbf{H}^{*}(10)$, but on the components of those measurands as they are provided by RadField3D, which are the volumetric photon fluence $\Phi^{3}_{\mathrm{\gamma}}$ and the volumetric photon energy distribution $p(E_{\mathrm{\gamma}})$ composed of $32$ bins and $E_{max}=150\,\mathrm{keV}$. In fact, the concurrent measurements of an area dosimeter are only needed to correct the fluence. Therefore, one can potentially use the estimated spectra to correct an area dosimeter measurement in turn.
@@
For the training process of our FCNNs, we used a cosine annealing [29] learning rate scheduler with $\mathrm{eta}_{min}=10^{-6}$ together with an Adam optimizer using weight decay [30] of $10^{-4}$ and $\beta_{1}=0.9\land\beta_{2}=0.99$. Additionally, we preceded our main learning rate scheduler by a linear warmup phase that reaches our target learning rate after $1000$ steps.
@@
Our network begins by feeding the location vector through a frequency encoding using sine and cosine functions with different periodicities which was initially introduced by A. Vaswani et al.[31] for the transformer networks and reused for the original NeRF implementation. Even though, the direction vector could also be represented by a simple frequency encoding, we instead used a spherical harmonics based encoding as the cyclic nature of an rotation angle can better be represented by that encoding. This was demonstrated for NeRF-like networks by InstantNGP of T. Müller et al.[32]. For efficiency reasons, we use the spherical harmonics implementation as it is provided by the tiny-cuda-nn package [33] with version 2.0. For both encodings, we append the raw input to the encoded vector. In contrast to the classical NeRF application, our network does not aim to learn the scatter objects 3D geometry with visual densities for volumetric rendering. Thus, we skip the pre-processing path where the model should optimize the voxel location representation towards the local density independently from the direction. We want the network to learn the implicit representation of the fluence distribution, which is similar to the visual density, but in our case is ultimately depended from the radiation direction. Thus, we extract the fluence as well as the spectrum from the last layer. In order build a robust backbone for solving the radiation field estimation under more complex parameter spaces, we do not just concatenate the feature vectors that describe the global beam conditions like the view direction in NeRF. Instead, we experimented with various feature fusing methods and finally compared the Feature-wise Linear Modulation (FiLM) [34] approach, which was originally used in CNNs, with a Gated Multimodal Unit (GMU) [35] approach and as the baseline, with the simple concatenation of feature vectors like it is done in classical NeRF networks. Ahead of fusing both feature vectors, the encoded X-ray tube direction vector is passed through the global encoding block, a Multilayer Perceptron (MLP) block that is similar to the MLP processing the encoded location vector with one hidden layer and a model width that is equal to the width of the main path of the network.
@@
The actual used hyperparameters for this and the extensions of this network, like frequencies of the frequency encoding, spherical harmonics degrees, model widths and feature fusing layers are reported in the results section as we determined them using hyperparameter tuning. Through all layers of the main path of this network, except for the last ones, we use SiLU activation functions. Only for the output activation we use different activation functions. For the fluence, we use a different activation function depending on the range of the fluence normalization. For the case of a normalization in $[0,1]$, we used a sigmoid $\sigma$ activation function, while for a normalization in [1,1][-1,1], we used a gradient conserving clamping $\mathrm{clamp}_{\mathrm{GC}}$, with:
@@
This network is the first extension of SRBFNet. It extends the global parameters g by the x-ray spectrum of the radiation source. Each radiation field file contains the $150$ bin histogram of the X-ray tube output spectrum using a bin width of $1\,\mathrm{keV}$. In order to improve the radiation field distribution and include the reconstruction of the local spectra, we encode that spectrum into our network. First, we reduce the dimensionality drastically by bilinear resampling the bins of the histogram down to $64$ bins. The reduced histogram is then passed through a MLP with one hidden layer using a width of $32$ neurons and the SiLU activation function. After the first linear layer, we implemented a layer-normalization to reshape the activation distribution.
The two feature vectors of the X-ray tube direction and the X-ray tube output spectrum are then concatenated and combined by the global encoding block in Fig. 5 that previously just processed the direction vector.
@@
This neural network targets the limitation of the pervious described variant, that is only able to learn a fixed, static beam distance. Therefore, this model adds an additional encoder to allow for the additional beam parameter introduced by DS-03 while keeping the same overall structure from SPERFNet. This requires one new scalar input parameter, that is the tube distance as the opening angle of the rectangular beam is directly depending on the distance. We observed, that simply passing this scalar raw is leading the neural network to ignore it. The previously used frequency encoding is not useful here either, as the input data is neither of high frequency nor periodical. Thus, we added a mini MLPs into the architecture, with has a width of $16$ neurons and in an optimal way encodes the distance information for the global parameters vector g.
@@
Instead of predicting every single voxel by itself, which is inefficient when the aim is to always predict all voxels, especially for fine resolutions, one can use CNNs. Those networks are relying on using convolution kernels, with multiple channels, that sweep over the image space and by that encoding local information. Initially, those networks were only used for Computer Vision, but by transposing this operation, those networks can be used for generative purposes.
@@
Our Beam2Scatter (B2S) U-Net is in general a CNN, that encodes the raw radiation field of the direct beam prior to the interaction with the patient geometry and from that should estimate the scatter radiation field as it is induced by the patient. Therefore, its encoder-decoder architecture follows the 3D-U-Net description [25], which itself is just an adaption to the third dimension of the original U-Net, as it was initially used for biomedical image segmentation [36]. We adopt this structure and give our network the fluence map of the direct beam as a one-channel input and let the network predict a $33$ dimensional output with the first $32$ channels for the spectra and the last channel for the fluence. Another addition is, that we introduced FiLM [34] layers after each up-sampling step in the decoder path, to modulate the weights of the decoder depending on the encoded input spectrum of the X-ray tube, which is encoded the same way as in SPERFNet. This way we can inject that information into the U-Net structure. The networks outputs got the prior introduced activation functions applied: ${\mathrm{clamp}_{\mathrm{GC}}}_{min}^{max}$ and $\mathrm{norm}_{\mathrm{hist}}$. Thanks to the direct beam radiation distribution being the input of this network, it can be directly used with all three datasets. For DS-01, the spectra encoding and fusing by using FiLM layers was disabled.
@@ 2.4 Training pipeline
To ensure that the models are all comparable, the training pipeline for all models was identical. Regardless, whether the model itself was capable of predicting partial or whole fields, the evaluation of the loss functions and the sub sequential update of the weights was performed in the same way. All models were trained for a maximum of $200$ epochs, which could be prematurely aborted, when the validation loss was not improving for the last $10$ epochs. For the training loop, we used the pytorch lightning [37] framework which implements the Cross-Validation policy with a final test using a dedicated validation and test dataset. The ratio of the split into training, validation and test dataset was $70\,\%$, $15\,\%$ and $15\,\%$. At first, a number of fields is loaded from the dataset that are transformed to pytorch tensors, preserving the channels and single layers, and as such they are batched together. We used a physical batch size of $4$ full radiation fields. Multiple batches of those predicted and ground-truth fields are aggregated together to an effective batch size of $64$ for training. We applied data pre-processings to each field prior to the prediction. Those pre-processings included the optional joining of the direct beam and scattered field channels, but also normalizing the ground-truth and input tensors. Finally, the loss functions for each of the neural networks normalized outputs (fluence and spectra) are calculated. Different from that, the metrics are calculated on the back-transformed data in the original data space.
The initial learning rates where estimated per neural network by using the pytorch lighting implementation of a learning rate finder, which follows the descriptions of L. N. Smith [38]. We limited it to search on the interval of $10^{-2},10^{-4}]$ for $250$ steps.
@@
For the FCNN models, which operate per voxel, we added an optional extra step to the training pipeline, that is typical for NeRF approaches. We refer to it as voxel location randomization, which is essentially a perturbation of the voxel position around its center with a maximum jitter of half the voxels dimension. This was used as a data augmentation to stabilize training in NeRF variants like InstantNGP [32].
@@
Across all our models we used the same two loss functions, that we optimize concurrently: $\pazocal{L}_{\Phi^{3}_{\mathrm{\gamma}}}$ the loss regarding the volumetric fluence $\Phi^{3}_{\mathrm{\gamma}}$ and $\pazocal{L}_{p(E_{\mathrm{\gamma}})}$ the loss regarding the local photon spectrum $p(E_{\mathrm{\gamma}})\in\mathbb{R}^{32}$ for $\sum_{i=0}^{31}p(E_{\mathrm{\gamma},i})=1$.
@@
For the hyperparameter tuning, we used the Optuna [39] framework with version 4.5. For the tuning of each network we defined the search space for possible parameters as listed in Tab. 2. The tuning was in general performed on every model we discuss: B2S U-Net, SRBFNet, SPERFNet, PBRFNet. Anyway, not all of the parameters could be tested for each model. For instance, the SRBFNet does not encode any spectra information, therefore the spectra encoding related parameters were not probed.
@@
normalizer $\mathrm{MaxNorm}_{0}^{1}$, $\mathrm{MaxNorm}_{-1}^{+1}$, $\mathrm{MaxLogNorm}_{10^{3}}$, $\mathrm{MaxLogNorm}_{10^{5}}$]

### v2

@@ 2.2 Data normalization
Figure 1: Distribution of relative voxel fluences inside datasets DS-01 and DS-02 for range $\Phi^{3}_{\mathrm{\gamma}}\geq 0.0$ on the X-axis, exhibiting a strong imbalance towards $\Phi^{3}_{\mathrm{\gamma}}\leq 10^{-2}$.
Figure 2: Distribution of relative voxel fluences inside datasets DS-01 and DS-02 for range $\Phi^{3}_{\mathrm{\gamma}}\geq 10^{-1}$ on the X-axis, exhibiting a much weaker imbalance in that range.
Figure 3: Distribution of relative voxel fluences inside dataset DS-03 for range $\Phi^{3}_{\mathrm{\gamma}}\geq 0.0$ on the X-axis, exhibiting a strong imbalance towards $\Phi^{3}_{\mathrm{\gamma}}\leq 10^{-5}$.
Figure 4: Distribution of relative voxel fluences inside dataset DS-03 for range $\Phi^{3}_{\mathrm{\gamma}}\geq 10^{-1}$ on the X-axis, exhibiting a strong imbalance towards $\Phi^{3}_{\mathrm{\gamma}}\leq 20^{-1}$ and a small peak at $\Phi^{3}_{\mathrm{\gamma}}=1$.
In principle, radiation fields inside the dataset are already normalized, as the photon spectra are normalized to probability distributions $p(E_{\mathrm{\gamma}})$ and the fluences are normalized per single primary photon from the simulation. However, due to the numerical nature of the datasets, after loading a single radiation field from a dataset, we performed an extra normalization step for the data to limit the occurring gradients and avoid numerical instability, as well as introducing bias because of high total values. Thus, this extra normalization step effectively mitigates the severe class imbalance in all our dataset towards low fluence voxels. This has been visualized in the global fluence histograms in Fig. 4 for DS-01 and DS-02 and in Fig. 4 for DS-03. The histograms were generated using the MaxNorm function to transform the fluences to $[0,1]$. From both figures we can draw that the majority of voxels have a fluence lower than $5\,\%$ of the maximum fluence, which makes sense, as the major part of the radiation field is indeed scattered radiation. However, this imbalance could potentially lead to a mis-reconstruction of the primary beam. Especially for DS-03 from Fig. 4, we can see that the class imbalance is also present inside the primary beam component. Not as severely expressed, datasets DS-01 and DS-02 exhibit an unequal distribution of fluences as well, as visible from Fig. 4. Therefore, we examine the influence of different normalization functions regarding the reconstruction capabilities, to find the best fit to reconstruct the high and low fluence regions with sufficient results.
@@
The idea of using alternative normalizations in the interval of $[0,1]$ is to create a more equal distribution of gradient strengths over the range of values to reconstruct. We hypothesize, that this is of importance for our use case, as the dosimetry is regarding relative instead of absolute errors per voxel. In general, it can be beneficial to elevate differences between small fluences and flatten the differences between higher fluences to give small deviations on small scales more impact. A classical solution for this issue is to use a transformation to the log-space, which is in general a reasonable choice for the following training of neural networks. Unfortunately, due to numerical instabilities, the steep ascent of the logarithm at very small values can become unstable under inversion and thus break the bijective property required for normalization. This instability can be observed for fluences of a magnitude lower than $10^{-8}$. As it can be seen from Tab. 1, for DS-03 fluences of about $10^{-8}$ play a role for the reconstruction as well. To mitigate this issue, we added a scaling factor $\alpha$ prior to the log-transformation. We tested two values for $\alpha$: $10^{0}$ and $10^{3}$. As an alternative to encounter the high dynamic range, we doubled the range of our previously linearly spread-out normalization from $\Phi^{3}_{\mathrm{\gamma}}\in[0,1]$ to $\Phi^{3}_{\mathrm{\gamma}}\in[-1,1]$ which additionally separates the low fluence voxels from the few high energy voxels by the sign. We distinguish between those two variants of the MaxNorm by indicating them as $\mathrm{MaxNorm}_{0}^{1}$ and $\mathrm{MaxNorm}_{-1}^{+1}$.
@@ 2.3 Architectures
In order to access the effectivity of our neural network architectures, we implemented each architecture’s neural network model in three variants to handle the different provided parameters for each dataset. With this approach, for each architecture there is one model that can only handle X-ray tube rotations, one that can additionally handle varying spectra and one that can additionally handle varying tube distances.
In general, our models all estimate the relative topology of the spatial radiation field distribution conserving the relation of intensities between all voxels inside one specific radiation field. This is important, as the actual fluence scales between different settings of the beam would introduce a bias towards high X-ray tube currents and thus harm the training of the neural networks. On the downside, in a real-world application, this inherently requires our approach to run concurrently to at least one area dosemeter that reports dose rates at a known point in space. As these devices, especially those with real-time measurement capabilities, are known to exhibit a certain energy and thus spectrum dependence, one needs to correct them accordingly. This is specifically of importance in our targeted use case, as we aim to predict the scattered field around the patient, where the spectrum is not homogeneous and is difficult to predict. Thus, we decided to train all our networks not using the actual target measurand, e.g., air kerma rate $\dot{K}_{\mathrm{air}}$ free-in-air or $H^{*}(10)$, but on the components of those measurands as provided by RadField3D, which are the volumetric photon fluence $\Phi^{3}_{\mathrm{\gamma}}$ and the volumetric photon energy distribution $p(E_{\mathrm{\gamma}})$ composed of $32$ bins and $E_{max}=150\,\mathrm{keV}$. In fact, the concurrent measurements of an area dosemeter are only needed to correct the fluence. Therefore, one can potentially use the estimated spectra to correct an area dosemeter measurement in turn.
@@
For the training process of our FCNNs, we used a cosine annealing [29] learning rate scheduler with $\mathrm{eta}_{\mathrm{min}}=10^{-6}$ together with an Adam optimizer using weight decay [30] of $10^{-4}$ and $\beta_{1}=0.9\land\beta_{2}=0.99$. Additionally, we preceded our main learning rate scheduler by a linear warmup phase that reaches our target learning rate after $1000$ steps.
@@
Our network begins by feeding the location vector through a encoder using frequency encoding using sine and cosine functions with different periodicities which were initially introduced by A. Vaswani et al.[31] for the transformer networks and reused for the original NeRF implementation. Even though, the direction vector can also be represented by a simple frequency encoding, we instead used a spherical harmonics based encoding as the cyclic nature of a rotation angle can be better represented by that encoding. This was demonstrated for NeRF-like networks by InstantNGP of T. Müller et al.[32]. For efficiency reasons, we used the spherical harmonics implementation as provided by the tiny-cuda-nn package [33] with version 2.0. For both encodings, we appended the raw input to the encoded vector. In contrast to the classical NeRF application, our network does not aim to learn the scatter objects 3D geometry with visual densities for volumetric rendering. Thus, we skipped the pre-processing path where the model should optimize the voxel location representation towards the local density independently from the direction. We want the network to learn the implicit representation of the fluence distribution, which is similar to the visual density, but in our case is ultimately dependent on the radiation direction. Thus, we extracted the fluence as well as the spectrum from the last layer. In order to build a robust backbone for solving the radiation field estimation under more complex parameter spaces, we did not just concatenate the feature vectors that describe the global beam conditions, such as the view direction in NeRF. Instead, we experimented with various feature fusing methods and finally compared the Feature-wise Linear Modulation (FiLM) [34] approach, which was originally used in CNNs, with a Gated Multimodal Unit (GMU) [35] approach and as the baseline, with the simple concatenation of feature vectors as is done in classical NeRF networks. Ahead of fusing both feature vectors, the encoded X-ray tube direction vector was passed through the global encoding block, a Multilayer Perceptron (MLP) block that is similar to the MLP processing the encoded location vector with one hidden layer and a model width that is equal to the width of the main path of the network.
@@
The actual hyperparameters used for this and the extensions of this network, such as the frequencies of the frequency encoding, spherical harmonics degrees, model widths and feature fusing layers are reported in the results section as we determined them using hyperparameter tuning. Through all layers of the main path of this network, except for the last ones, we use SiLU activation functions. Only in the case of the output activation did we use different activation functions. For the fluence, we used a different activation function depending on the range of the fluence normalization. For the case of a normalization in $[0,1]$, we used a sigmoid $\sigma$ activation function, while for a normalization in [1,1][-1,1], we used a gradient conserving clamping $\mathrm{clamp}_{\mathrm{GC}}$, with:
@@
This network is the first extension of SRBFNet. It extends the global parameters g by the x-ray spectrum of the radiation source. Each radiation field file contains the $150$ bin histogram of the X-ray tube output spectrum using a bin width of $1\,\mathrm{keV}$. In order to improve the radiation field distribution and include the reconstruction of the local spectra, we encoded that spectrum into our network. First, we reduced the dimensionality drastically by bilinear resampling the bins of the histogram down to $64$ bins. The reduced histogram was then passed through a MLP with one hidden layer using a width of $32$ neurons and the SiLU activation function. After the first linear layer, we implemented a layer-normalization to reshape the activation distribution.
The two feature vectors of the X-ray tube direction and the X-ray tube output spectrum were then concatenated and combined by the global encoding block in Fig. 5 that previously processed the direction vector.
@@
This neural network targets the limitation of the previous described variant, that is only able to learn a fixed, static beam distance. Therefore, this model adds an additional encoder to allow for the additional beam parameter introduced by DS-03 while keeping the same overall structure from SPERFNet. This requires one new scalar input parameter, that is the tube distance as the opening angle of the rectangular beam is directly dependent on the distance. We observed, that simply passing this scalar raw leads the neural network to ignore it. The previously used frequency encoding is not useful here either, as the input data is neither of high frequency nor periodical. Thus, we added a mini MLPs into the architecture, which has a width of $16$ neurons and in an optimal way encodes the distance information for the global parameters vector g.
@@
Instead of predicting every single voxel by itself, which is inefficient when the aim is to always predict all voxels, especially for fine resolutions, one can use CNNs. These networks rely on using convolution kernels with multiple channels that sweep over the image space and by that, encodes local information. Initially, these networks were only used for Computer Vision, but by transposing this operation, these networks can be used for generative purposes.
@@
Our Beam2Scatter (B2S) U-Net is in general a CNN, that encodes the raw radiation field of the direct beam prior to the interaction with the patient geometry and from that estimates the scattered radiation field as it is induced by the patient. Therefore, its encoder-decoder architecture follows the 3D-U-Net description [25], which itself is just an adaption to the third dimension of the original U-Net, as it was initially used for biomedical image segmentation [36]. We adopted this structure and gave our network the fluence map of the direct beam as a one-channel input and let the network predict a $33$ dimensional output with the first $32$ channels for the spectra and the last channel for the fluence. Another addition is, that we introduced FiLM [34] layers after each up-sampling step in the decoder path, to modulate the weights of the decoder depending on the encoded input spectrum of the X-ray tube, which is encoded the same way as in SPERFNet. This way we can inject that information into the U-Net structure. The networks outputs had the prior introduced activation functions applied: ${\mathrm{clamp}_{\mathrm{GC}}}_{min}^{max}$ and $\mathrm{norm}_{\mathrm{hist}}$. Thanks to the direct beam radiation distribution being the input of this network, it can be used directly with all three datasets. For DS-01, the spectra encoding and fusing by using FiLM layers was disabled.
@@ 2.4 Training pipeline
To ensure that the models are all comparable, the training pipeline for all models was identical. Regardless, whether the model itself was capable of predicting partial or whole fields, the evaluation of the loss functions and the sub sequential update of the weights was performed in the same way. All models were trained for a maximum of $200$ epochs, which could be prematurely aborted when the validation loss was not improving for the last $10$ epochs. For the training loop, we used the pytorch lightning [37] framework which implements the Cross-Validation policy with a final test using a dedicated validation and test dataset. The ratio of the split into training, validation and test dataset was $70\,\%$, $15\,\%$ and $15\,\%$. At first, a number of fields were loaded from the dataset that are transformed to pytorch tensors, preserving the channels and single layers, and as such they were batched together. We used a physical batch size of $4$ full radiation fields. Multiple batches of those predicted and ground-truth fields were aggregated together to an effective batch size of $64$ for training. We applied data pre-processing to each field prior to the prediction. This pre-processing included not only the option of joining the direct beam and scattered field channels, but also normalizing the ground-truth and input tensors. Finally, the loss functions for each of the neural networks normalized outputs (fluence and spectra) are calculated. The metrics are calculated on the inverse-transformed data in the original data space.
The initial learning rates were estimated per neural network by using the pytorch lighting implementation of a learning rate finder, which follows the descriptions of L. N. Smith [38]. We limited it to search on the interval of $10^{-2},10^{-4}]$ for $250$ steps.
@@
For the FCNN models, which operate per voxel, we added an optional extra step to the training pipeline, that is typical for NeRF approaches. We refer to it as voxel location randomization, which is essentially a perturbation of the voxel position around its center with a maximum jitter of half the voxel’s dimension. This was used as a data augmentation to stabilize training in NeRF variants like InstantNGP [32].
@@
Across all our models we used the same two loss functions, that we optimized concurrently: $\pazocal{L}_{\Phi^{3}_{\mathrm{\gamma}}}$ the loss regarding the volumetric fluence $\Phi^{3}_{\mathrm{\gamma}}$ and $\pazocal{L}_{p(E_{\mathrm{\gamma}})}$ the loss regarding the local photon spectrum $p(E_{\mathrm{\gamma}})\in\mathbb{R}^{32}$ for $\sum_{i=0}^{31}p(E_{\mathrm{\gamma},i})=1$.
@@
For the hyperparameter tuning, we used the Optuna [39] framework with version 4.5. For the tuning of each network, we defined the search space for possible parameters as listed in Tab. 2. The tuning was, in general, performed on every model we discuss: B2S U-Net, SRBFNet, SPERFNet, PBRFNet. Anyway, not all of the parameters could be tested for each model. For instance, the SRBFNet does not encode any spectra information, therefore the spectra encoding related parameters were not probed.
@@
normalizer $\mathrm{MaxNorm}_{0}^{1}$, $\mathrm{MaxNorm}_{-1}^{+1}$, $\mathrm{MaxLogNorm}_{10^{0}}$, $\mathrm{MaxLogNorm}_{10^{3}}$]